HETAV PATEL

Grade 11,

what is resistance across 1 side of a 9 ohm equilateral triangle

what is resistance across 1 side of a 9 ohm equilateral triangle
 

Grade:12

2 Answers

Khimraj
6 years ago
Two resistance will be in series and their equivalent will be parallel to other so equivalent is = (9+9)9/{(9+9)+9} = 6 ohm.Hope it clears. If u have doubts then please clearify. And if u like my answer then please approve answer
hero
6 years ago
If each resistance of side os 9 ohm then two resistance are in series and their resultant is parallel to one side . Resultant of series is 
9+9 that is 18 ohm which is connected to parallel resistance of 9 ohm hence resultant is 1/R=1/9+1/18
THAT IS R=6 OHM